2. The percent composition of a compound was found to be 63.5% silver, 8.2% nitrogen, and 28.3 oxygen. Determine the compounds empirical formula. A 170.00 g sample of an unidentified compound contains 29.84 g sodium, 67.49 chromium, and 72.67 g oxygen. What is the compounds empirical formula? A tetraethyl lead, a gasoline additive, is found to contain 38.43 g 60.00 g sample of hydrogen. Find its empirical formula. lead, 17.83 g carbon, and 3.74 g
Answer #1

Q2.

MW of Silver = 107.8682 g/mol

MW of Nitrogen = 14 g/mol

MW of Oxygen = 16 g/mol

so..

assume a basis of 100 g

so

63.5 g of Silver, 8.2 g of N, 28.3 g of O

mol of Ag = mass/MW = 63.5/107.8682 =0.588681 mol

mol of N= mass/MW = 8.2/14= 0.5857 mol

mol of O= mass/MW = 28.3/16=1.76875 mol

ratio is:

1.76875/0.588681 = 3

Ag:N:3O

Empiricla formula:

AgNO3

i.e.

Silver Nitrate